The angular momentum of an electron is J, its kinetic energy will be?

Grade:11

3 Answers

Swayam Singh
6 years ago
according to Bohr
angular momentum of an electron = nh/2(pi)
hence,
mvr = nh/2(pi)
according to question angular momentum is “J”
means 
nh/2(pi) = J
now 
mvr = nh/2(pi)
squaring both sides
m2v2r2 = n2h2/4(pi)2
m2v2r= J2
now knietic energy = 1/2mv2
hence,
multiplying both sides my ½
1/2m2v2r2 = J2/2
1/2mv2 = J2/2r2m
Akanksha
6 years ago
We knoe that Mvr = n h/ 2πMvr= JMv = J/r Squaring both sides M2v2 =J2/R2Mv2=j2/R2 *mAs we know kinetic energy is 1/2mv2Ke= 1/2 mv2 =1/2 j2/r2 m
Akanksha
6 years ago
We knoe that Mvr = n h/ 2πMvr= JMv = J/r Squaring both sides M2v2 =J2/R2Mv2=j2/R2 *mAs we know kinetic energy is 1/2mv2Ke= 1/2 mv2 =1/2 j2/r2 m